OpenStudy (anonymous):

the slope of a line is -4 and its y-intercept is (0,4). what is the equation of the line that is perpendicular to the first line and passes through (-8,2)

OpenStudy (anonymous):

general linear equation y=mx+c m=gradient/slope c=y-intercept we are given gradient=-4 y-intercept = 4 y=-4x + 4 <<this is the equation we can come up with. perpendicular meaning we have to follow the rule m1 x m2 = -1 m1=gradient of our known equation which we figured above (-4) m2=gradient of our unknown equation which we want to find. -4 x m2 = -1 m2 = -1/-4 m2 = 1/4 use the point (-8,2) and m=1/4 to substitute into the general linear equation to find ur c value then re-write the whole thing.